# Waveform Design of Multi-User-Multi-Target ISAC System based on Kullback–Leibler Divergence

### v1

@@ Abstract
This paper presents a novel framework that leverages the Kullback-Leibler divergence (KLD) to analyze and optimize performance trade-offs in integrated sensing and communication (ISAC) systems. We consider a multiple-input-multiple-output (MIMO) base station that simultaneously serves communication user equipments (UEs) and detects multiple targets using shared antenna deployment. The proposed KLD-based approach provides a unified performance measure encompassing both UE error rate and target detection capability. We apply this approach on two well-known communication beamforming techniques, maximum ratio transmission (MRT) and zero-forcing (ZF), and evaluate the effect on the radar subsystem. Furthermore, two optimization problems are formulated and solved. The first one optimizes the KLD of the radar subsystem for given constraints on the communication KLD, whereas the second one focuses on communication waveform KLD-based optimization and constrained radar KLD. These optimization problems are solved using a projected gradient method with an adaptive penalty for the radar waveform and a gradient-assisted interior point method for the communication waveform. Through theoretical derivations and extensive simulations, it is demonstrated that our approach can be a powerful tool for characterizing and optimizing the performance trade-offs of ISAC under various configurations. The results also show significant improvements in both sensing and communication performance by the KLD-optimized system compared to well-known benchmarks, such as conventional MRT and ZF for the communication subsystem, and the conventional identity covariance design for the radar subsystem. These findings support the holistic design and optimization of ISAC in next-generation wireless networks.
@@
Integrated sensing and communication, multiple-input-multiple-output (MIMO), radar, zero-forcing, maximum ratio transmission (MRT), beamforming, optimization, Kullback–Leibler divergence (KLD).
@@ I Introduction
The rapid evolution of wireless communication networks, particularly with the advent of fifth-generation (5G) technology, has dramatically expanded the landscape of interconnected devices and services. This technological leap has enabled a wide array of applications that depend heavily on advanced sensing capabilities. From autonomous vehicles and unmanned aerial vehicles (UAVs) to the Internet-of-Things (IoT) ecosystems, these applications are pushing the boundaries of traditional communication systems, necessitating the integration of sophisticated sensing functionalities [1, 2, 3, 4]. As we stand on the threshold of the sixth-generation (6G) era, network operators and service providers face unprecedented challenges in not only delivering robust communication services but also in integrating advanced sensing-based applications such as detection, localization, tracking, navigation, and environmental surveillance [5, 6, 7].
Integrated Sensing and Communications (ISAC) systems have garnered significant interest in the research and industrial communities to meet these emerging demands. ISAC systems aim to synergistically merge sensing and communication functionalities within the same infrastructure, thereby optimizing the use of available resources at base stations (BSs) for dual purposes [8, 9, 10, 11, 12, 13]. This integration can be achieved through various architectural approaches. For example, in a separated deployment, the BS antennas are allocated distinctly between communication and radar subsystems, which can simplify the system design but may underutilize resources. In contrast, the shared deployment strategy employs all antennas for both functionalities, potentially offering higher resource efficiency at the cost of the increased complexity and the need for more sophisticated signal processing techniques[14, 15].
The evolution of ISAC systems has been accelerated by the advances in multi-antenna technologies, particularly massive Multiple-Input Multiple-Output (MIMO) systems. These systems allow for the generation of highly directional beams and support simultaneous service to a large number of users, thereby significantly enhancing both communication capacity and sensing accuracy, usually done by exploiting spatial diversity by setting the antenna separation to a minimum of half wavelength [16]. The synergy between ISAC and massive MIMO presents a promising avenue for meeting the stringent requirements of future wireless networks. However, evaluating the performance of ISAC systems poses unique challenges due to their dual functionality. Traditionally, communication and sensing subsystems have been assessed using distinct metrics, with communication performance measured by achievable rate, outage probability, and bit error rate, while the sensing performance is typically evaluated using metrics like estimation rate, detection probability, false alarm probability, and mean square error (MSE) [17]. This disparity complicates the holistic assessment and optimization of ISAC systems.
@@ I-A Literature review
To address these challenges mentioned above, there has been a growing interest in developing unified performance measures that can simultaneously capture both sensing and communication capabilities. One such measure that has gained traction is the Kullback-Leibler divergence (KLD), also known as the relative entropy. KLD has a well-established history in information theory and has been extensively used in the analysis and design of sensing systems, particularly in the context of MIMO radars [18, 19, 20]. It quantifies the dissimilarity between two probability distributions, serving as a measure of the information gain achieved by using one distribution over another. While the application of KLD in sensing systems is well-established, its potential in characterizing communication system performance remains open. Recent studies have begun to bridge this gap, demonstrating the utility of KLD in assessing the performance trade-offs between sensing and communication functionalities in ISAC systems. For instance, in our previous works [21, 22], we employed KLD to analyze performance in systems with separated antenna deployment, while in [23] we proposed a low-complexity unified objective function based on KLD, aimed at optimizing network resources for both separated and shared deployment scenarios. Moreover, further insights about the achievable KLD trade-off in ISAC systems is explored in [24], and more designs have been discovered. Nonetheless, the literature lacks a comprehensive investigation and beamforming design for a generalized multi-user-multi-target (MUMT) ISAC system using the promising KLD.
@@
Further efforts have explored various optimization approaches for ISAC systems, each addressing specific aspects of system performance. For example, joint radar-communication beamforming optimization based on the Cramér-Rao Bound (CRB) has been proposed to improve both sensing and communication performance [25]. This approach minimizes the CRB of angle estimation for radar while maximizing the signal-to-interference-plus-noise ratio (SINR) for communication, thus enhancing the accuracy of sensing and the quality of communication simultaneously. However, these methods often focus on optimizing individual components rather than adopting a unified framework that captures the interplay between different functionalities. Beamforming optimization has been a particular focus, with studies investigating partially connected hybrid beamforming designs to address the hardware complexity of fully digital systems [26]. While these approaches offer practical advantages in terms of implementation, they do not fully capture the nuanced trade-offs between sensing and communication performance, especially in scenarios where these functionalities are deeply integrated. Waveform optimization has also been extensively studied, with recent work proposing a symbol-level precoding scheme using a faster-than-Nyquist approach to enhance the spectral efficiency in bandwidth-constrained environments [27]. This approach focuses on communication waveform design while considering the radar functionality as a constraint, ensuring that the designed waveforms maintain certain radar performance metrics. However, this work does not consider a unified optimization framework that could simultaneously address the needs of both radar and communication subsystems.
@@ I-B Motivation and contribution
Despite these advances, current approaches often fall short of fully exploiting the potential of integrated systems. Most notably, there is a need for more comprehensive models that can accurately reflect the real-world operation of ISAC systems, particularly in scenarios where communication signals can be utilized for radar target detection alongside dedicated radar signals. Exploiting the interference from communication signals as a source of information for radar operations enhances radar detection capabilities and maximises the system’s overall resource efficiency [14]. This dual-use of communication signals represents a significant departure from traditional approaches, which typically treat interference as a detrimental factor to be mitigated rather than leveraged[28, 29].
Toward these objectives and motivated by the fact that KLD is able to provide a unified measure for ISAC system holistically, this paper proposes a novel framework based on the Kullback-Leibler divergence (KLD) to characterize and optimize the sensing-communication performance trade-offs in shared antenna deployment ISAC systems. This work addresses several gaps in the existing literature by introducing a unified KLD-based performance measure that captures the synergies and trade-offs between sensing and communication in ISAC systems. By considering the shared deployment scenario and proposing specific optimization techniques for both radar and communication waveforms, we provide a comprehensive approach for system optimization that reflects the true potential of ISAC systems in future wireless networks. Additionally, our analysis of different beamforming techniques within this framework offers valuable guidance for designing ISAC systems that can effectively balance the needs of both radar and communication functionalities. Our system model consists of a multiple-input-multiple-output (MIMO) base station (BS) that simultaneously serves communication user equipment (UEs) and detects multiple targets using shared resources. The key contributions of our work are as follows:
@@
  • Inherent Exploitation of Communication Interference: Our framework accounts for the fact that communication signals can also be exploited for radar target detection alongside dedicated radar signals, providing a more comprehensive and useful model of ISAC system operation. This dual-use approach enhances the radar performance by utilizing communication interference as an additional source of information for target detection.
@@
  • KLD-based Optimized ISAC: We propose two novel optimization techniques: radar and communication waveform KLD-based optimization. Both maximize their respective KLD measures subject to the minimum KLD of both subsystems and power constraints. The non-convex radar optimization employs a projected gradient method with an adaptive penalty, while the communication optimization uses a gradient-assisted interior point method. These techniques efficiently handle ISAC waveform design’s complex, multi-dimensional nature, enabling balanced performance optimization across both subsystems within system constraints.
@@
  • Benchmark Analysis: We investigate two benchmark beamforming techniques for the communication subsystem, namely maximum ratio transmission (MRT) and zero-forcing (ZF), and analyze their impact on the overall system performance within our KLD-based framework. This analysis provides practical insights for system designers seeking to optimize ISAC performance in shared deployment scenarios with a complexity-performance trade-off.
@@
Our results demonstrate that the proposed KLD-based optimization techniques significantly enhance ISAC system performance compared to the MRT and ZF scenarios for the communication subsystem and the radar subsystem’s conventional identity covariance design. The radar waveform optimization shows substantial improvements in both target detection and communication performance, while the communication waveform optimization primarily benefits the communication subsystem with modest radar performance gains. Both techniques exhibit robust performance across varying SNR levels, with the radar optimization demonstrating particularly stable computational efficiency. Notably, incorporating communication signals in radar detection significantly enhances overall system performance, especially in scenarios with limited dedicated radar resources. These findings provide valuable insights for designing efficient shared deployment ISAC systems that balance sensing and communication requirements.
I-C Paper organization
@@
The paper is structured as follows. Section II presents the system model. Sections III and IV analyze the communication and radar systems, respectively, including the derivations of KLD for each system. In Section V, the first KLD-based optimisation technique considers the radar KLD sum under communication KLD constraint is investigated. Section VI introduces the second KLD-based optimisation technique based on communication KLD sum with radar KLD constraint. Section VII displays the numerical results, and finally, Section VIII concludes the work.
@@ II System Model
This paper is concerned in an ISAC system that comprises an $N$ antenna MIMO-BS. The antennas are utilised for detecting a maximum number of ${T}$ targets and serving $K$ number of single-antenna communication user equipments (UEs) in the downlink direction using the shared deployment scenario. The total transmitted power available at BS is $P_{\mathrm{T}}$ which is utilised for both sensing and data communication duties. The portions of the power that are allocated to the radar and communication subsystems are respectively denoted as $P_{\mathrm{r}}$ and $P_{\mathrm{c}}$, where $P_{\mathrm{T}}=P_{\mathrm{c}}+P_{\mathrm{r}}$. As a starting point, non-optimised beamforming techniques such as MRT and ZF are employed at BS to precode the information of communication UEs [16]. These techniques serve as benchmarks for our subsequent optimized systems, which will be introduced and analyzed later in Sections V and VI. The combined transmitted ISAC signal $\mathbf{x}_{l}\in\mathbb{C}^{N\times 1}$ will be as follows,
@@
where $\mathbf{s}_{\mathrm{c},l}\in\mathbb{C}^{K\times 1}$ is a vector of communication UEs symbols, $\mathbf{s}_{\mathrm{r}}\in\mathbb{C}^{T\times 1}$ is a vector of the baseband radar waveforms for the potential targets, $\mathbf{W}_{\mathrm{c}}\in\mathbb{C}^{N\times K}$ represents the precoding matrix for the communication subsystem, and $\mathbf{W}_{\mathrm{r},l}\in\mathbb{C}^{N\times T}$ represents the precoding matrix for the radar subsystem. Notably, the transmitted signals vector $\mathbf{x}_{l}$ is transmitted from the BS and received at the UEs as well as the targets which reflect the received version of $\mathbf{x}_{l}$ back to the BS. Although the radar signal component received at the UEs might impose interference that cannot be eliminated by the users, the communication signal component received at the targets and reflected back to the BS carries information about the target. Therefore, by carefully designing the precoding matrices at the BS, the sensing information can be enhanced by exploiting the communication signal for sensing targets as well. This dual-use approach enhances radar performance by utilizing the communication signal as an additional source of information. Moreover, the radar waveform design in our work is based on the combined transmit signal covariance matrix for the $t$-th beam such that the covariance matrix of the $t$-th transmit beam is $\mathbf{R}_{t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{x}_{t,l}\>\mathbf{x}_{t,l}^{H}$, where $\mathbf{x}_{t,l}=\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}+\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}$, and $L$ is the total number of snapshots. This formulation allows us to incorporate the communication signals into the radar detection process. It should be noticed that the power for the radar is integrated into the precoding vector $\mathbf{w}_{\mathrm{r},t,l}$, where it is designed to have $\|\mathbf{w}_{\mathrm{r},t,l}\|_{F}^{2}=\frac{P_{\mathrm{r}}}{T}$, and $\mathbf{w}_{\mathrm{r},t,l}=\frac{P_{\mathrm{r}}}{NT}\>\mathbf{\tilde{w}}_{\mathrm{r},t,l}$, where $\mathbf{\tilde{w}}_{\mathrm{r},t,l}$ is the $t$th beam precoding vector, and .$\|.\|_{F}$ is the Frobenius norm.

### v2

# A Framework for Holistic KLD-based Waveform Design for Multi-User-Multi-Target ISAC Systems

@@ Abstract
This paper introduces a novel framework that leverages the Kullback-Leibler Divergence (KLD) metric to analyse and optimise performance trade-offs in integrated sensing and communication (ISAC) systems. We consider a multiple-input multiple-output (MIMO) base station that simultaneously serves communication user equipments (UEs) and detects multiple targets using a shared antenna deployment. We apply this framework to two widely used communication beamforming techniques, maximum ratio transmission (MRT) and zero-forcing (ZF), to assess their impact on the radar subsystem’s performance. Additionally, two optimisation problems are formulated: the first optimises the radar subsystem’s KLD under communication constraints, and the second focuses on communication waveform KLD optimisation with constraints on the radar KLD. These problems are solved using a projected gradient method with adaptive penalties for the radar waveforms and a gradient-assisted interior point method (IPM) for the communication waveforms. Through theoretical derivations and extensive simulations, we demonstrate that our KLD approach effectively characterises and optimises the performance trade-offs between sensing and communication in ISAC systems. The results show significant improvements in both radar detection and communication performance when compared to traditional MRT and ZF beamforming, and the identity covariance design for radar subsystems. These findings promote a more holistic design and optimisation of ISAC for next-generation wireless networks and demonstrate the advantages of KLD-based optimisation in balancing the performance of both sensing and communication.
@@
Integrated sensing and communication, multiple-input-multiple-output (MIMO), radar, zero-forcing, maximum ratio transmission, beamforming, optimisation, Kullback–Leibler divergence (KLD).
@@ I Introduction
The rapid evolution of wireless communication networks has dramatically expanded the landscape of interconnected devices to include new applications such as autonomous vehicles and unmanned aerial vehicles (UAVs) that are heavily dependent on advanced sensing, pushing the boundaries of communication systems beyond their traditional limits [1, 2, 3, 4]. As we approach the sixth-generation (6G) era, network operators and service providers aim not only to further enhance their communication services but also to include advanced sensing-based applications such as detection, localisation, tracking, navigation, and environmental surveillance [5, 6, 7].
Integrated Sensing and Communications (ISAC) systems have gained significant interest as a key enabler of many 6G applications. ISAC aims to synergistically merge sensing and communication functionalities hence optimising base station (BS) resources for dual purposes [8, 9, 10, 11, 12, 13]. This integration can be achieved through various approaches. For example, in a separated deployment, the BS antennas are allocated distinctly between communication and radar subsystems, which can simplify the system design but may underutilise resources. In contrast, the shared deployment strategy employs all antennas for both functionalities, hence offering higher resource efficiency, but at the cost of increased complexity due to the need for more sophisticated signal processing techniques[14, 15].
The evolution of ISAC has been accelerated by advances in multi-antenna technologies, particularly massive Multiple-Input Multiple-Output (MIMO) systems. These systems allow for the generation of highly directional beams and support simultaneous services to a large number of users, thereby significantly enhancing both communication capacity and sensing accuracy [16]. The synergy between ISAC and massive MIMO presents a promising avenue for meeting the stringent requirements of future wireless networks. However, evaluating the performance of ISAC systems poses unique challenges due to their dual functionality. Traditionally, communication and sensing subsystems have been assessed using distinct metrics, with communication performance measured by achievable rate, outage probability, and bit error rate (BER), while the sensing performance is typically evaluated using metrics like estimation rate, detection probability, false alarm probability, and mean square error (MSE) [17]. This disparity complicates the holistic assessment and optimisation of ISAC systems.
@@ I-A Literature review
To address these challenges mentioned above, there has been a growing interest in developing unified performance measures that can simultaneously capture both sensing and communication capabilities. One such measure that has gained traction is the Kullback-Leibler divergence (KLD), also known as the relative entropy. KLD has a well-established history in information theory and has been extensively used in the analysis and design of sensing systems, particularly in the context of MIMO radars [18, 19, 20]. It quantifies the dissimilarity between two probability distributions, serving as a measure of the information gain achieved by using one distribution over another. While the application of KLD in sensing systems is well-established, its potential in characterising communication system performance remains open. Recent studies have begun to bridge this gap, demonstrating the utility of KLD in assessing the performance trade-offs between sensing and communication functionalities in ISAC systems. For instance, in our previous works [21, 22], we employed KLD to analyse performance in systems with separated antenna deployment, while in [23] we proposed a low-complexity unified objective function based on KLD, aimed at optimising network resources for both separated and shared deployment scenarios. Moreover, further insights about the achievable KLD trade-off in ISAC systems are explored in [24], and more designs have been discovered. Nonetheless, the literature lacks a comprehensive investigation and beamforming design for a generalised multi-user-multi-target (MUMT) ISAC system using the promising KLD.
@@
Further efforts have explored various optimisation approaches for ISAC systems, each addressing specific aspects of system performance. For example, joint radar-communication beamforming optimisation based on the Cramér-Rao Bound (CRB) has been proposed to improve both sensing and communication performance [25]. This approach minimises the CRB of angle estimation for radar while maximising the signal-to-interference-plus-noise ratio (SINR) for communication, thus enhancing the accuracy of sensing and the quality of communication simultaneously. However, these methods often focus on optimising individual components rather than adopting a unified framework that captures the interplay between different functionalities. Beamforming optimisation has been a particular focus, with studies investigating partially connected hybrid beamforming designs to address the hardware complexity of fully digital systems [26]. While these approaches offer practical advantages in terms of implementation, they do not fully capture the nuanced trade-offs between sensing and communication performance, especially in deeply integrated scenarios. Waveform optimisation has also been extensively studied, with recent work proposing a symbol-level precoding scheme using a faster-than-Nyquist approach to enhance the spectral efficiency in bandwidth-constrained environments [27]. This approach focuses on communication waveform design while considering the radar functionality as a constraint, ensuring that the designed waveforms maintain certain radar performance metrics. However, this work does not consider a unified optimisation framework that could simultaneously address the needs of both subsystems.
@@ I-B Motivation and contribution
Despite these advances, current approaches often fall short of fully exploiting the potential of integrated systems. Most notably, there is a need for more comprehensive models that can accurately reflect the real-world operation of ISAC systems, particularly in scenarios where communication signals can be utilised for radar target detection alongside dedicated radar signals. Exploiting the interference from communication signals as a source of information for radar operations enhances radar detection capabilities, improving the system’s overall resource efficiency [14]. This dual-use of communication signals represents a significant departure from traditional approaches, which typically treat interference as a detrimental factor to be mitigated rather than leveraged[28, 29].
Toward these objectives and motivated by the fact that KLD is able to provide a unified measure for ISAC system holistically, this paper proposes a novel framework based on the Kullback-Leibler divergence (KLD) to characterise and optimise the sensing-communication performance trade-offs in shared antenna deployment ISAC systems. This work addresses several gaps in the existing literature by introducing a unified KLD-based performance measure that captures the synergies and trade-offs between sensing and communication in ISAC systems. By considering the shared deployment scenario and proposing specific optimisation techniques for both subsystem waveforms, we provide a comprehensive approach for system optimisation that reflects the true potential of ISAC systems in future wireless networks. Additionally, our analysis of different beamforming techniques within this framework offers valuable guidance for designing ISAC systems to effectively balance the needs of both subsystem functionalities. Our system model consists of a multiple-input-multiple-output (MIMO) base station (BS) that simultaneously serves communication user equipment (UEs) and detects multiple targets using shared resources. The key contributions of this work are:
@@
  • Exploitation of Communication Interference: Our framework accounts for the fact that communication signals can also be exploited for radar target detection alongside dedicated radar signals, providing a more comprehensive and useful model of ISAC system operation. This dual-use approach enhances the radar performance by utilising communication interference as an additional source of information for target detection.
@@
  • KLD-based Optimised ISAC: We propose two novel optimisation techniques: radar and communication waveform KLD-based optimisation. Both maximise their respective KLD measures subject to the minimum KLD of both subsystems and power constraints. The non-convex radar optimisation employs a projected gradient method with an adaptive penalty, while the communication optimisation uses a gradient-assisted interior point method. These techniques efficiently handle ISAC waveform design’s complex, multi-dimensional nature, enabling balanced performance optimisation across both subsystems within system constraints.
@@
  • Benchmark Analysis: We investigate two benchmark beamforming techniques for the communication subsystem, namely maximum ratio transmission (MRT) and zero-forcing (ZF), and analyse their impact on the overall system performance within our KLD-based framework. This analysis provides practical insights for optimising ISAC performance in shared deployment scenarios, considering the complexity-performance trade-off.
@@
Our results demonstrate that the proposed KLD-based optimisation techniques significantly enhance ISAC system performance compared to the MRT and ZF scenarios for the communication subsystem and the radar subsystem’s conventional identity covariance design. The radar waveform optimisation shows substantial improvements in both target detection and communication performance, while the communication waveform optimisation primarily benefits the communication subsystem with modest radar performance gains. Both techniques exhibit robust performance across varying SNR levels, with the radar optimisation demonstrating particularly stable computational efficiency. Notably, incorporating communication signals in radar detection significantly enhances overall system performance, especially in scenarios with limited dedicated radar resources. These findings provide valuable insights for designing efficient shared deployment ISAC systems that balance sensing and communication requirements.
I-C Paper organisation
@@
The paper is structured as follows. Section II presents the system model. Sections III and IV analyse the communication and radar systems, respectively, including the derivations of KLD for each system. In Section V, the first KLD-based optimisation technique considers the radar KLD sum under communication KLD constraint is investigated. Section VI introduces the second KLD-based optimisation technique based on communication KLD sum with radar KLD constraint. Section VII displays the numerical results, and finally, Section VIII concludes the work.
@@ II System Model
We consider an $N$ antenna MIMO-BS. The antennas are utilised for detecting a maximum number of ${T}$ targets and serving $K$ number of single-antenna communication UEs in the downlink direction in shared deployment manner. The total transmitted power available at BS is $P_{\mathrm{T}}$ which is utilised for both sensing and data communication duties. The power $P_{\mathrm{T}}$ allocated to the radar and communication subsystems are denoted as $P_{\mathrm{r}}$ and $P_{\mathrm{c}}$, where $P_{\mathrm{T}}=P_{\mathrm{c}}+P_{\mathrm{r}}$. As a starting point, non-optimised beamforming techniques such as MRT and ZF are employed at BS to precode the information of communication UEs [16]. These techniques serve as benchmarks for our subsequent optimised systems, which will be introduced and analysed later in Sections V and VI. The combined transmitted ISAC signal $\mathbf{x}_{l}\in\mathbb{C}^{N\times 1}$ will be as follows,
@@
where $\mathbf{s}_{\mathrm{c},l}\in\mathbb{C}^{K\times 1}$ is a vector of communication UEs symbols, $\mathbf{s}_{\mathrm{r}}\in\mathbb{C}^{T\times 1}$ is a vector of the baseband radar waveforms for the potential targets, $\mathbf{W}_{\mathrm{c}}\in\mathbb{C}^{N\times K}$ represents the precoding matrix for the communication subsystem, and $\mathbf{W}_{\mathrm{r},l}\in\mathbb{C}^{N\times T}$ represents the precoding matrix for the radar subsystem. The transmitted signals vector $\mathbf{x}_{l}$ is transmitted from the BS and received at the UEs as well as the targets which reflect the received $\mathbf{x}_{l}$ back to the BS. Although the radar signal component received at the UEs might cause interference that cannot be eliminated by the users, the communication signal component received at the targets and reflected back to the BS carries information about the target. Therefore, by carefully designing the precoding matrices at the BS, the sensing information can be enhanced by exploiting the communication signal. This dual-use approach enhances the radar performance by utilising the communication signal as an additional source of information. The radar waveform design is based on the combined transmit signal covariance matrix for the $t$-th beam where the covariance matrix of the $t$-th transmit beam is $\mathbf{R}_{t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{x}_{t,l}\>\mathbf{x}_{t,l}^{H}$, $\mathbf{x}_{t,l}=\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}+\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}$, and $L$ is the total number of snapshots. This formulation allows us to incorporate the communication signals into the radar detection process. It should be noticed that the power for the radar is integrated into the precoding vector $\mathbf{w}_{\mathrm{r},t,l}$, where it is designed to have $\|\mathbf{w}_{\mathrm{r},t,l}\|_{F}^{2}=\frac{P_{\mathrm{r}}}{T}$, and $\mathbf{w}_{\mathrm{r},t,l}=\frac{P_{\mathrm{r}}}{NT}\>\mathbf{\tilde{w}}_{\mathrm{r},t,l}$, where $\mathbf{\tilde{w}}_{\mathrm{r},t,l}$ is the $t$th beam precoding vector, and .$\|.\|_{F}$ is the Frobenius norm.